Calculate the pH of a solution with a hydroxide concentration of 0.0112.
1
Expert's answer
2020-05-28T10:33:25-0400

pOH = - lg(OH-) = - lg(0.0112) = 1.95

pH = 14 - pOH = 14 - 1.95 = 12.05